If a cylinder has a volume of 339 cubic inches and a diameter of 6 inches, what is the height of the cylinder.
Answer by rfer(10417)

V=pir^2h
h=V/pir^2
h=339/3.14(3)^2
h=339/3.14*9
h=339/28.26
h=12 in